OpenStudy (hap1236):

Find the exact value of the expression. Do not use a calculator. If tan θ = 9, find the exact value of cot (0) A. 8 B. 10 C.1/9 D. 9

OpenStudy (hap1236):

someone please help

zepdrix (zepdrix):

\[\large\rm \cot \theta=\frac{1}{\tan \theta}\]Flip the number

OpenStudy (hap1236):

i dont know how to do it

OpenStudy (hap1236):

1/9

TheSmartOne (thesmartone):

that's correct, :) and I'm assuming the question meant to say cot (θ) instead of cot (0) θ in place of 0 :P